### v2

# RedOne: Revealing Domain-specific LLM Post-Training in Social Networking Services

Abstract
As a primary medium for modern information dissemination, social networking services (SNS) have experienced rapid growth, which has proposed significant challenges for platform content management and interaction quality improvement. Recently, the development of large language models (LLMs) has offered potential solutions but existing studies focus on isolated tasks, which not only encounter diminishing benefit from the data scaling within individual scenarios but also fail to flexibly adapt to diverse real-world context. To address these challenges, we introduce RedOne, a domain-specific LLM designed to break the performance bottleneck of single-task baselines and establish a comprehensive foundation for the SNS. RedOne was developed through a three-stage training strategy consisting of continue pretraining, supervised fine-tuning, and preference optimization, using a large-scale real-world dataset. Through extensive experiments, RedOne maintains strong general capabilities, and achieves an average improvement up to 14.02% across 8 major SNS tasks and 7.56% in SNS bilingual evaluation benchmark, compared with base models. Furthermore, through online testing, RedOne reduced the exposure rate in harmful content detection by 11.23% and improved the click page rate in post-view search by 14.95% compared with single-tasks finetuned baseline models. These results establish RedOne as a robust domain-specific LLM for SNS, demonstrating excellent generalization across various tasks and promising applicability in real-world scenarios.
RedOne: Revealing Domain-specific LLM Post-Training
in Social Networking Services
Fei Zhao, Chonggang Lu, Yue Wang, Zheyong Xie, Ziyan Liu, Haofu Qian, JianZhao Huang, Fangcheng Shi, Zijie Meng, Hongcheng Guo, Mingqian He, Xinze Lyu, Yiming Lu, Ziyang Xiang, Zheyu Ye, Chengqiang Lu, Zhe Xu, Yi Wu, Yao Hu, Yan Gao, Jun Fan, Xiaolong Jiang, Weiting Liu, Boyang Wang, Shaosheng Caothanks: Corresponding author. NLP Team, Xiaohongshu Inc., China caoshaosheng@xiaohongshu.com
1 Introduction
With the widespread adoption of online platforms and mobile applications, social networking services (SNS) have emerged as a central medium for modern information dissemination, such as communication, knowledge sharing, and emotional expression Elahimanesh et al. (2025). Unlike the general textual corpora, SNS data is highly informal, context-sensitive, and often emotionally charged. These characteristics present unique challenges including linguistic variability, frequent role-switching, and subtle conversational norms, which complicate applications (e.g. platform content management and interaction quality improvement) for traditional natural language processing (NLP) systems Jin et al. (2024).
Figure 1: Performance comparison of different models in the SNS domain, where all models are instruction-tuned and the evaluation score is the average of all tasks on SNS-Bench.
Given these complexities, numerous studies have explored recent advanced large language models (LLMs) based adaptation for SNS-related tasks Zeng et al. (2024); Jiang and Ferrara (2023). However, these solutions primarily focus on isolated tasks, which not only experience diminishing benefits as data scales within individual scenarios but also struggle to adapt flexibly to diverse real-world contexts. This highlights a fundamental limitation in current SNS domain-specific models, where performance plateaus due to the inability to incorporate a more diverse domain knowledge corpus during training Yue et al. (2025).
To address these deficiencies, we introduce RedOne, a demain-specific LLM with a meticulous three-stage post-training strategy using a large-scale dataset from real-world, which consists of continued pretraining (CPT), supervised fine-tuning (SFT), and preference optimization (PO). In the CPT stage, the model acquire extensive foundational knowledge in the SNS domain by processing large-scale corpora. Building on this foundation, the SFT stage refines the model’s capability to tackle specific SNS tasks by leveraging carefully defined domain-specific problem formulations. Finally, in the PO stage, we further optimize the model’s behavior to ensure seamless alignment with human preferences and maximize its practical utility in real-wold deployments.
Through extensive experiments, RedOne not only maintains strong general capabilities, but also excels across multiple SNS-specific evaluation benchmarks, significantly outperforming leading proprietary or open-source models as shown in Figure 1. Further online testing in harmful content detection and post-view search, indicates its broad and promising potential application in real-world scenarios.
Our contributions can be summarized as follows:
  • We introduce RedOne, a domain-specific LLM, engineered to break the performance bottleneck of single-task models, providing comprehensive improvements for SNS.
  • A three-stage training strategy is designed, using a large-scale real-world dataset, which maintains strong general capabilities while delivering exceptional generalization across diverse SNS tasks.
  • Through extensive experiments and online testing to demonstrate RedOne’s effectiveness across a wide range of tasks, and establish a comprehensive and robust baseline for SNS application.
2 Related Work
2.1 NLP tasks in Social Networking Services
Due to the inherent characteristics of SNS platforms, namely their informality and rapid linguistic evolution Carr and Hayes (2015), these platforms present numerous complex NLP challenges that have garnered sustained academic attention. In the early stages of development, researchers primarily focused on fundamental capability assessments, particularly prevalent tasks such as sentiment analysis Mohammad et al. (2018); Rosenthal et al. (2019), harmful content detection i Orts (2019); Lu et al. (2024), and meme detection Xie et al. (2023); Lin et al. (2024). Following the emergence of LLMs and building upon previous research foundations, various techniques have evolved in multiple domains, including content understanding Kumar et al. (2024); Kmainasi et al. (2024), information extraction Islam and Goldwasser (2025); Li et al. (2024b); Peng et al. (2024), and dialogue systems Yi et al. (2024); Zhang et al. (2024). These technological advances have significantly enhanced problem-solving capabilities within the SNS domain, but have primarily focused on single tasks. In contrast to these works, RedOne demonstrates superior performance across diverse SNS tasks, providing a foundational model for improved services.
2.2 Domain-specific Post-training
To better serve specialized domains, recent efforts have focused on developing vertical domain LLMs across various fields, including finance Wu et al. (2023); Konstantinidis et al. (2024), law Colombo et al. (2024), home renovation Wen et al. (2023), medicine Xiong et al. (2023); Chen et al. (2023); Yang et al. (2024c); Wu et al. (2024); Zakka et al. (2024), and scientific research Azerbayev et al. (2023); Bi et al. (2023); Yang et al. (2024d). Despite these advancements, these vertical domain LLMs have not addressed the unique challenges posed by SNS. While Liu et al. (2024b) and Yang et al. (2024b) explore the application of LLMs to a limited set of NLP tasks within SNS, their coverage remains constrained. Therefore, a significant gap exists in this area, which RedOne aims to address.
3 RedOne Model
Figure 2: Overview of our training pipeline.
As illustrated in Figure 2, the training strategy of RedOne contains three stages. First, in Section 3.1, we conduct continue pretraining to enrich the model’s grasp of nuanced SNS field knowledge. Subsequently, in Section 3.2, we sharpen the model’s instruction-following capabilities through supervised fine-tuning across various tasks. Finally, we leverage preference information from the training data to perform preference optimization, ultimately yielding the RedOne with superior performance in the SNS domain.
3.1 Continue Pretraining
To enhance the large model’s fundamental domain knowledge we conducted continue pretraining at this stage, which can be divided into three sub-stages: data collection and data construction, filtering and mixture, along with domain-aware continue pretraining.
3.1.1 Data Collection and Data Construction
We specifically collected continue pretraining data from the following two data sources:
(1) General High-quality Data. We selected several high-quality open-source pretraining corpora Qiu et al. (2024); Weber et al. (2024); Penedo et al. (2024) to preserve the model’s fundamental generalization capabilities. To improve training efficiency, we uniformly construct all general data into single-sentence text format and perform segmentation and concatenation processing based on predefined text length thresholds.
(2) SNS-specific Domain Data. We collect the large-scale training data from SNS platforms and the open web, capturing diverse social communication patterns including informal discussions, short-form comments, sarcasm, emotionally charged content, and so on. For better reveal the underlying information in the pretraining data, we incorporate user interaction data to guide the training process. Specifically, we group contexts and comments with their corresponding user interaction data, which naturally clusters semantically related SNS content without additional processing. Through these steps, we collected and constructed a large-scale dataset comprising various tasks with over 100B tokens for downstream processing.
3.1.2 Filtering and Mixture
Considering data quality is crucial for model training Zhou et al. (2023a), we constructed a data-filtering pipeline inspired by Yuan et al. (2024), which comprises task-oriented rule filtering and small language model filtering Wang et al. (2025). The former identifies specific error content such as HTML tags and repetitive sentences, while the latter focuses on global assessment aspects including coherence and tone appropriateness. Based on this data-filtering pipeline, we further applied the RegMix method Liu et al. (2024a) to identify an optimal data mixture distribution and filter out unnecessary data. Through this comprehensive filtering and mixture process, we ultimately constructed a high-quality dataset of 20B tokens for training.
3.1.3 Domain-aware Continue Pretraining
After data construction, we conduct continue pretraining on the complete dataset. Specifically, RedOne is trained from the Qwen2.5 Qwen et al. (2025a) checkpoint using the same configurations, leveraging its strong linguistic capabilities across multiple domains. Through this domain-aware continue pretraining process, we ultimately obtain a model that effectively captures SNS-specific linguistic patterns while maintaining minimal degradation in general language modeling capabilities.
3.2 Supervised Fine-Tuning
To bridge the gap between pretraining objectives and the specific requirements of real-world SNS applications, we further conduct supervised fine-tuning on our model through carefully designed data construction and multi-stage training strategies.
3.2.1 Task Definition and Data Construction
As SFT training data is significat affect the final instruction following ability in domain tasks Dong et al. (2023), we extensively collected large-scale user-generated content from public platforms, including notes, comments, queries, and interaction logs, which provide real enviorment signal for us to improve model actions. Notably, we focused on preserving the linguistic style which exhibit typical SNS characteristics such as informal language, sarcasm, sentiment, and topical shifts while collecting data Eisenstein (2013), aim for representative and practical coverage for SNS scenarios.
After data collection, we ultimatly consolidate six kinds of core capabilities essential for SNS applications: content understanding, information extraction, semantic matching, user behavior modeling, dialogue and persona simulation, and translation, as show in Table 1. Each is supported by well-defined tasks reflecting real-world challenges and the overview is shown in Appendix A.2.
Additionally, during SFT, we also incorporated open source instruction data covering general tasks such as instruction following Li et al. (2025); Zhou et al. (2023a), multiturn dialogue Zhao et al. (2024), and long chain-of-thought (CoT) reasoning Guha et al. (2025); Ye et al. (2025), to mitigate catastrophic forgetting McCloskey and Cohen (1989) and retain generalization ability of RedOne model.
Task Name Capability
Note Taxonomy Content Understanding
Query Classification Content Understanding
Query Intent Recognition Content Understanding
Hashtag Prediction Information Extraction
Machine Reading Comprehension Information Extraction
Highlight Word Detection Information Extraction
Query-Note Relevance Semantic Matching
Query-Note Retrieval Semantic Matching
Post-View Search User Behavior Modeling
Emotional Companion Dialogue Dialogue
Role-playing Dialogue Dialogue
SNS Domain Translation Translation
Table 1: Overview of SNS Tasks and Their Capabilities
3.2.2 Two-Step Training
In domain SFT, a two-step mixed fine-tuning has been demonstrated to effectively enhance domain-specific capabilities Dong et al. (2024). For RedOne’s SFT, we implement this strategy by mixing SNS-specific data with general data across two steps. In the first step, we train the model on the complete SNS dataset combined with a large volume of general data. This approach enables the model to learn diverse task formats within the SNS domain while preserving its generalization capabilities. In the second step, we fine-tune the model using a higher proportion of SNS domain data, thereby further enhancing performance on domain-critical tasks.
3.3 Preference Optimization
SNS tasks like query-note relevance modeling often produce multiple plausible but quality-diverse outputs. While SFT improves instruction-following, it fails to exploit implicit preference signals among these candidates, causing overfitting and poor generalization Chu et al. (2025). To address these limitations, in this section, we carefully craft preference data and perform PO to obtain a better domain-specific model.
3.3.1 Preference Data Construction
To enhance alignment with human preferences and utilize the information embedded in data labels, we integrate different preference pair construction strategies according to the nature of different task types. Specifically, we categorize our data into two types and adopt corresponding strategies:
For subjective tasks, such as emotional dialogue and role-playing, our primary objective is to achieve better alignment with human preferences. Therefore, the first step begins with domain experts creating preference annotations on model-generated responses Ouyang et al. (2022). Furthermore, to scale up the preference dataset, we evaluate the consistency between trained judge models Cao et al. (2024a) and human preference, then leverage these models with high performance to expand specific data.
In contrast, for objective tasks with definitive correct answers, our strategy shifts toward extracting and utilizing the implicit structural information within the data labels. Here, we employ two approaches: First, we leverage the inherent structure of questions that contain both correct answers and incorrect options, constructing preference pairs that exploit the ordinal relationships within data. Complementarily, to actively address model limitations, we construct preference pairs from model errors, using ground truth as positive examples and incorrect predictions as negative to target specific weaknesses.
By integrating these tailored approaches, we systematically process all SNS-domain data according to their inherent characteristics, ultimately constructing preference optimization datasets that effectively capture both human preferences and implicit data information for comprehensive model enhancement.
3.3.2 Direct Preference Optimization
To effectively leverage the rich preference signals in our SNS dataset, we adopt DPO Rafailov et al. (2023) as our preference-based fine-tuning algorithm. This approach enables the model to better align with human preferences while simultaneously exploiting the latent information embedded in ground-truth labels.
Finally, through this comprehensive three-stage training pipeline encompassing CPT, SFT and PO, we ultimately obtain a domain-specific large language model RedOne that demonstrates superior performance in the target domain while maintaining reasonable general capabilities.
4 Experiments
4.1 Implementation details
During the CPT stage, we follow the training process from Qwen2.5 (Yang et al., 2024a) over a mixed corpus of general and SNS-specific data. SFT is conducted for three epochs in step one and two epochs in step two, with a maximum sequence length of 16 38416\,384 using sequence packing, batch size of $128$, a linear warm-up ratio of 0.1. The learning rates are set according to model size: for the 7B model, we use $5\text{\times}{10}^{-6}$ in step one and $3\text{\times}{10}^{-6}$ in step two; for the 32B model, we use $3\text{\times}{10}^{-6}$ for both steps. Optimization is performed using AdamW Loshchilov and Hutter (2017) ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\epsilon{=}10^{-8}$). In the final PO stage, we employ a learning rate of $1\times 10^{-7}$, batch size of $64$, sequence length of $4096$, training for two epochs, with SFT loss coefficient set to 0.3.
Models General-Bench SNS-Bench SNS-TransBench
Avg. Taxonomy Hashtag QueryCorr MRC NER Gender CHLW QueryGen Avg. ZH→EN EN→ZH Avg.
BLEU chrF++ BLEU chrF++
Llama-3.1-8B Grattafiori et al. (2024) 51.24 37.74 66.62 33.32 31.27 47.10 74.61 26.88 38.60 44.52 23.07 48.15 29.32 29.13 32.42
Ministral-8B Mistral-AI (2024) 49.93 42.62 70.58 36.24 30.71 37.79 82.38 28.04 46.27 46.83 25.67 50.91 32.02 31.18 34.95
InternLM3-8B Cai et al. (2024) 58.55 51.83 76.98 38.65 25.25 39.41 66.84 44.71 43.46 48.39 24.85 50.44 35.58 34.04 36.23
GLM-4-9B-0414 GLM et al. (2024) 63.27 56.03 77.67 38.03 45.29 47.01 51.30 27.51 45.52 48.55 32.20 56.90 39.73 37.40 41.57
Qwen2.5-7B Qwen et al. (2025b) 63.01 49.50 73.80 42.37 45.32 45.41 88.08 33.76 44.65 52.86 31.43 55.91 38.36 36.48 40.55
RedOne-7B (Ours) 63.83 (+0.82%) 72.18 88.02 65.09 63.98 51.86 70.47 74.73 48.69 66.88 (+14.02%) 38.06 62.66 46.88 44.82 48.11 (+7.56%)
Table 2: Results of 7B-scale models. Bold entries indicate the best model, while underlined entries denote the second one. Percentage improvements relative to the baseline Qwen2.5 foundation model are also shown.
Models General-Bench SNS-Bench SNS-TransBench
Avg. Taxonomy Hashtag QueryCorr MRC NER Gender CHLW QueryGen Avg. ZH→EN EN→ZH Avg.
BLEU chrF++ BLEU chrF++
Open-Source Large Language Models
Phi-4-14B Abdin et al. (2024) 63.00 57.62 79.56 46.32 53.39 44.99 89.12 29.23 44.76 55.62 31.28 57.23 37.58 36.68 40.69
Mistral-Small-24B Mistral-AI (2025) 65.63 64.88 83.89 48.77 46.51 52.09 91.19 32.10 46.01 58.18 31.29 56.72 39.28 37.32 41.15
Llama-3.3-70B Grattafiori et al. (2024) 67.64 62.94 83.28 50.76 27.38 56.09 91.19 33.58 46.41 56.45 34.00 59.18 41.25 39.56 43.50
GLM-4-32B-0414 GLM et al. (2024) 74.39 63.36 85.50 47.33 53.72 50.41 80.31 33.19 46.90 57.59 36.32 61.31 42.53 40.77 45.23
Deepseek-V3-0324 DeepSeek-AI et al. (2025) 75.22 67.27 86.59 47.71 60.97 56.00 90.16 40.45 46.03 61.90 35.65 61.58 46.86 44.58 47.17
Closed-Source Large Language Models
Doubao-1.5-Pro-32k Doubao-Team (2025) 76.13 30.00 83.21 58.25 61.32 56.60 90.67 30.61 46.55 57.15 33.71 61.85 45.54 44.35 46.36
GLM-4-Plus GLM et al. (2024) 70.25 65.46 84.31 52.13 55.81 53.16 86.53 30.09 44.68 59.02 41.57 65.95 48.79 47.06 50.84
GPT-4o-1120 OpenAI 70.72 65.79 84.98 51.79 58.89 54.99 88.08 38.96 47.33 61.35 40.32 63.91 49.15 47.28 50.17
Claude-3.7-Sonnet Anthropic 75.10 72.03 88.83 54.10 54.86 56.13 92.23 31.11 45.49 61.85 35.63 61.66 45.79 44.23 46.83
Gemini-2.0-Flash DeepMind (2024) 74.42 68.76 87.36 48.41 52.21 53.58 89.64 37.39 46.27 60.45 32.72 58.84 41.80 40.16 43.38
Qwen-Max Qwen et al. (2025b) 71.86 65.68 84.47 54.36 61.34 55.78 91.19 37.97 46.64 62.18 35.55 60.92 46.08 44.14 46.67
Qwen2.5-32B Qwen et al. (2025b) 71.68 59.90 80.51 46.00 55.04 54.51 90.67 38.84 45.66 58.89 32.56 58.14 42.34 40.71 43.44
RedOne-32B (Ours) 73.72 (+2.04%) 81.45 90.19 67.07 59.24 51.66 81.87 70.40 50.37 69.03 (+10.14%) 40.55 64.54 48.20 46.05 49.84 (+6.40%)
Table 3: Results of 32B-scale models. Bold entries indicate the best model, while underlined entries denote the second one. Percentage improvements relative to the baseline Qwen2.5 foundation model are also shown.
4.2 Benchmarks
For general capabilities evaluation, we use datasets similar to those employed in community, including general natural language comprehension (i.e. MMLU Hendrycks et al. (2021a), CMMLU Li et al. (2023a), CEVAL Huang et al. (2023), GPQA-Diamond Rein et al. (2023), NewsBench Li et al. (2024a)), reasoning (i.e. MMLU-Pro Wang et al. (2024), BBH Suzgun et al. (2023), GaokaoBench Zhang et al. (2023)), mathematics (i.e. AIME2025 MAA (2025), GSM8K Cobbe et al. (2021) and MATH500 Hendrycks et al. (2021b)), coding (i.e. HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and LiveCodeBench(24072502)Jain et al. (2024)), translation (i.e. WMT-22/23/24 and FloresGoyal et al. (2022)), instruction following (i.e. IFEval Zhou et al. (2023b)), hallucination and human preference alignment (i.e. HaluEval Li et al. (2023b) and CompassBench Cao et al. (2024b)). To further evaluate RedOne’s performance in SNS domain, we selected specialized SNS benchmarks including SNS-Bench Guo et al. and SNS-TransBench Guo et al. (2025).
4.3 Main Results
Figure 3: Model capability radar diagram across different task categories.
As shown in Tables 2 and 3, we conducted a comparison between RedOne with baseline models across various tasks in three categories. Meanwhile, as illustrated in Figure 3, we compared RedOne-7B and RedOne-32B with their base model across seven dimensions (six for general and one for SNS). Both results indicate that RedOne in all scales not only maintain robust general capabilities, even surpassing their base model on general tasks, but also exhibit exceptional effectiveness in the SNS domain. Additionally, RedOne achieves performance comparable to significantly larger models across most tasks, with limited improvement opportunities observed only in few areas. The results also demonstrate that scaling up RedOne consistently enhances performance over smaller variants, aligning with established model scaling laws. These findings underscore RedOne’s strong potential for further advances through continued increases in model size, as well as its promise for real world application.
4.4 Task-specific SFT Comparison
To further explore the impact of base model selection on task-specific fine-tuning and validate our domain LLM’s effectiveness, we conducted experiments on two 7B-scale models: the original Qwen-2.5-Instruct (“Qwen”) and our SNS-adapted model (“RedOne”). We evaluated three variants: (1) Qwen-Finetuned, involving task-specific fine-tuning on Qwen; (2) RedOne-Finetuned, involving task-specific fine-tuning on RedOne; and (3) RedOne, representing zero-shot inference without further fine-tuning.
Models HashTag QueryCorr MRC
Qwen2.5-Finetuned 88.93 57.76 62.26
RedOne 88.02 (-0.91%) 65.09 (+12.63%) 63.98 (+2.76%)
RedOne-Finetuned 90.51(+1.78%) 65.77(+13.87%) 64.47(+3.55%)
Models CHLW QueryGen SNS-Trans
Qwen2.5-Finetuned 78.41 48.25 48.01
RedOne 74.73 (-4.72%) 48.69 (+0.91%) 48.11 (+0.21%)
RedOne-Finetuned 79.11(+0.89%) 49.21(+1.99%) 48.32(+0.65%)
Table 4: Performance comparison of task-specific Finetuned on Qwen-2.5-Instruct and RedOne (all models are 7B scale).
As shown in Table 4, RedOne-Finetuned consistently outperforms Qwen2.5-Finetuned across most datasets, demonstrating that domain-aligned post-training (i.e., RedOne) provides a stronger foundation for downstream SFT. Meanwhile, even RedOne in the zero-shot setting exhibits strong performance, further corroborating the benefits of domain adaptation. Overall, these results indicate that initializing SFT from a domain-adapted base model is more effective than starting from a general-purpose large model. This finding suggests that domain-specific post-training can serve as a powerful approach for improving both zero-shot capabilities and task-specific performance after fine-tuning.
4.5 Ablation Study
We also investigate the contributions of each stage in our training pipeline, with results summarized in Table 5. While CPT alone shows limited immediate gains, it establishes a crucial foundation for subsequent specialization. Based on the CPT model, adding SFT and PO brings average improvements of 0.55 and 1.90 on SNS-Bench compared to variants without CPT, indicating that CPT provides a stronger knowledge foundation for SFT to improve instruction-following and also offers a broader exploration space for PO. Additionally, both CPT and SFT lead to a performance drop on general benchmarks, whereas PO could effectively mitigate this decline and further enhance the overall results. Finally, the complete three-stage pipeline yields the strongest results on specialized benchmarks, with 66.88 on SNS and 48.11 on SNS-Trans, while maintaining competitive general-domain performance at 63.83, demonstrating the effectiveness of our training strategy.
CPT SFT PO General SNS SNS-Trans
63.01 52.86 40.55
$\checkmark$ 62.65 64.57 47.47
$\checkmark$ $\checkmark$ 64.36 64.98 47.64
$\checkmark$ 62.28 53.28 41.39
$\checkmark$ $\checkmark$ 61.95 65.12 47.70
$\checkmark$ $\checkmark$ $\checkmark$ 63.83 66.88 48.11
Table 5: Ablation study results of RedOne-7B.
4.6 Online Results
Task Metric Change (%)
Harmful Content Detection Exposure Rate ($\downarrow$) 11.23-11.23
Post-View Search Click Page Rate ($\uparrow$) +14.95+14.95
Table 6: Effectiveness in online scenarios.
To further validate RedOne’s practical effectiveness, we deployed the model across multiple internal SNS scenarios and witnessed remarkable performance gains in real-world applications compared with previous single-task models as shown in Table 6. In harmful content detection, RedOne exhibited exceptional safety capabilities by slashing the exposure rate of harmful notes by 11.23%, effectively filtering out non-compliant content and strengthening platform security. Moreover, for post-view search recommendation, the model delivered a 14.95% increase in click page rate, indicating improved content discovery and enhanced user engagement following note interactions. These online results demonstrate the strong practical utility of RedOne in real-world scenarios.
4.7 Out-of-Domain Ability Analysis
Figure 4: Performance on OOD tasks for models of varying parameter size.
In this subsection, we examine the impact of preserving general-domain capabilities during domain adaptation by evaluating out-of-domain (OOD) robustness. Specifically, we select one task without corresponding supervised training data (Note Taxonomy) and two tasks with available data (Note Hashtag, Note MRC) from the SNS bench. For the latter two, we remove their related training data during supervised fine-tuning (SFT), effectively making all three tasks OOD.
We then compare models trained with both general and SNS data against those trained with SNS data only, across various model sizes. Our results (Figure 4) suggest that including general-domain data helps models generalize better to OOD tasks, and this trend is more visible for larger models. This highlights that maintaining general capabilities can be beneficial for domain adaptation, though further study is needed to fully understand this effect.
4.8 Case Study
Input Title: Found it! A Soft-Sole Commuter Loafer You Can Walk In All Day.
Tags: Height Increasing Thick Sole Shoes, Loafers, Beautiful Loafers.
Popular Comments: How to buy; link.
Qwen How to choose height-increasing platform shoes.
RedOne height-increasing thick-soled loafers.
Table 7: Post-view search case: Input SNS context and model-generated queries.
To further demonstrate RedOne’s effectiveness in capturing user search intent within SNS scenarios, we present a case study on the post-view search task. As shown in Table 7, we analyze a sample post featuring height-increasing loafers that generated significant purchase intent among users. Qwen produces a general shopping query, whereas RedOne directly identifies the core product keywords, better reflecting users’ intent to search for and purchase the featured item. This demonstrates RedOne’s superior capability for generating actionable queries aligned with real user needs.
5 Conclusion
In this paper, we introduce RedOne, a domain-specific LLM trained through a three-stage strategy that enhances SNS-specific capabilities while preserving general performance. We believe our approach can inspire future research in developing specialized LLMs and advancing practical applications in social media.
Limitations
Although our proposed method demonstrates strong effectiveness, it requires extensive data processing, resulting in considerable resource costs. Additionally, current models are still in a large scale, which increases online inference latency and serving expenses, limiting deployment in resource-constrained settings. Future work will explore lighter architectures, including model compression through quantization and distillation, as well as routing-efficient designs such as mixture-of-experts (MoE), to reduce latency and cost without sacrificing accuracy.
Ethical Considerations
When integrating large language models as essential components within application services, it is crucial to rigorously consider potential model hallucinations and security risks. To ensure reliable service delivery, leveraging RedOne for model services requires careful implementation to mitigate adverse user impacts. Furthermore, we emphasize the critical importance of adhering to stringent user privacy protection standards throughout data collection and processing workflows, ensuring comprehensive personal information security.
Reproducibility
Due to data privacy and code security concerns, we are currently unable to fully release our datasets and code. However, our entire development pipeline is built upon widely adopted open-source projects and works Zheng et al. (2024); Rafailov et al. (2023); Dong et al. (2023), which ensures that users only need to organize their data according to the specified format in order to run it smoothly. In addition, we have reported detailed parameter settings used during training in Section 4.1, which further provides valuable references for the community to reproduce our pipeline.
Appendix A Appendices
A.1 SNS Tasks
In this section, we provide an overview of the key tasks defined for SFT in SNS scenarios. These tasks are designed to reflect real-world user behavior, content patterns throughout social platforms. The SFT task suite covers six core capability areas, each capturing an important aspect of SNS applications:
Content Understanding:
This category focuses on the model’s ability to comprehend and categorize user-generated content as well as user queries. Example tasks include classifying notes into categories (note taxonomy), determining the topic or domain of user queries (query classification), and identifying fine-grained query intent (query intent recognition).
Information Extraction:
Tasks in this category address the identification and extraction of structured information from informal SNS posts. This includes predicting appropriate hashtags for a post, answering questions about note content, and detecting highlight or anchor words that represent user focus.
Semantic Matching:
Here, the model is required to judge the semantic relationship and relevance between items such as user queries and social notes. Typical tasks include evaluating whether a note is relevant to a given query (query-note relevance) and retrieving the most pertinent or high-quality notes for search scenarios (query-note retrieval).
User Behavior Modeling:
This capability involves modeling and simulating user actions, such as generating follow-up queries based on previous browsing or posting activities (post-view search). It reflects how users might interact with content in a dynamic SNS environment.
Dialogue and Persona Simulation:
To enhance natural interaction and personalization, dialogue tasks ask the model to engage in emotional companion conversations or role-play as different personas in group chats, capturing both the style and richness of real SNS dialogues.
Translation:
Given the prevalence of multilingual content, the model is also trained to translate notes between languages, with attention to preserving the original tone, sentiment, and informal expressions common across SNS platforms.
Each task adopts the most suitable instruction-tuning format: multiple choice supports classification and selection, extraction is used for entity and span prediction, and generation handles open-ended responses such as dialogue or translation. This format-driven design ensures consistent prompting and facilitates efficient multi-task training.
Figure 5: Token length distribution in the dataset. The histogram uses a logarithmic y-axis with dashed lines indicating the median (345 tokens) and the 95-th percentile (2,342 tokens).
(a) Primary task categories.
(b) Secondary task categories.
Figure 6: Top 10 distributions of primary and secondary task categories in the SFT dataset.
A.2 SFT Data Statistical Analysis
Token Length Statistics
Figure 5 presents the token length distribution across all samples in our dataset, displayed on a logarithmic scale to accommodate the wide range of sequence lengths up to 16,384 tokens. The distribution exhibits the characteristic heavy-tailed pattern typical of natural language corpora.
Task Category Distribution
We adopted the labeling taxonomy from Infinity Instruct Li et al. (2025), organizing it into primary and secondary categories. Specifically, we used a subset of Infinity Instruct data, consisting of instructions paired with their corresponding labels, as training data to fine-tune a labeling model based on Qwen2.5-7B-Instruct. This trained labeling model was then applied to annotate all instructions in our complete SFT dataset. The comprehensive distribution of primary and secondary label categories is illustrated in Figures 6(a) and 6(b), respectively.
The distribution analysis reveals several key characteristics of our SFT dataset. At the primary level, natural language processing and understanding dominates with over 3.4 million instances, followed closely by information processing and integration (3.1 million) and problem solving and support (2.2 million). This indicates a strong emphasis on core language comprehension and analytical capabilities. Mathematical ability and programming-related tasks also constitute significant portions, with over 1.2 million instances each, reflecting the dataset’s comprehensive coverage of technical skills. At the secondary level, information extraction leads with 2.1 million instances, while text understanding and problem-solving tasks follow with 1.7 million and 1.4 million instances respectively. The relatively balanced distribution across different cognitive abilities suggests that our dataset provides diverse training scenarios for developing well-rounded AI capabilities.